OpenStudy (anonymous):

i will be a fan of who helps me with this: Prove: sin θ - sin θ•cos2 θ = sin3 θ. You must show all work.

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ \sin θ - \sin θ\times\cos^2 θ = \sin3 θ }\) \(\normalsize\color{blue}{ \sin θ (1-\cos^2 θ )= \sin3 θ }\) divide by sin (on both sides, and apply `sin²x+cos²x=1`

OpenStudy (solomonzelman):

Tell me if you need more help

OpenStudy (bradely):

sinθ - sin θ•cos2 θ sinθ(1-cos^2θ) sinθ*sin^2θ =sin^3θ

OpenStudy (anonymous):

awesome! i will and thanks :D both of you

OpenStudy (solomonzelman):

Anytime

OpenStudy (solomonzelman):

I was assuming, that it is \(\normalsize\color{blue}{ \sin θ (1-\cos^2 θ )= \sin^3 θ }\)

OpenStudy (solomonzelman):

we can do \(\normalsize\color{blue}{ \sin θ (1-\cos(2θ)~~~ )= \sin(3θ) }\) if you need :)

OpenStudy (solomonzelman):

And btw, \(\normalsize\color{blue}{ \sin θ (1-\cos(2θ)~~~ )= \sin(3θ) }\) 1) Click and hold ALT 2) click the number code (using the numbers that are on the right of the keyboard, not the ones below F1, F2, F3, etc., ) 3) release the ALT `0 1 8 5` ¹ `2 5 3` ² `0 1 7 9` ³ `0 2 1 5` × `2 4 6` ÷ `7 5 4` ≥ `7 5 5 ` ≤ `2 4 1` or `7 5 3` ± `2 4 7` ≈ `2 5 1` √

OpenStudy (solomonzelman):

just some alt codes if you want... have a good day !

OpenStudy (anonymous):

thanks :) you too!

OpenStudy (solomonzelman):

Yw